# On the Domination of surface-group representations in $\mathrm{PU}(2,1)$

### v1

@@ 2.3. The $T$-bent representations and their ${\tt Z}$-invariants
This part mainly revisits [Wil12] and is included here for the sake of completeness. We begin by recalling a geometric fact about pairs of adjacent real ideal triangles, which underlies the definition of the ${\tt Z}$-invariant.
@@
$\tau_{0}=\bigg{(}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}-1\\ -\sqrt{2}\\ 1\end{bmatrix},\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}\bigg{)}\ \text{ and }\ \tau_{z}=\bigg{(}\begin{bmatrix}1\\ 0\\ 0\end{bmatrix},\begin{bmatrix}0\\ 0\\ 1\end{bmatrix},\begin{bmatrix}-|z|^{2}\\ z\sqrt{2}\\ 1\end{bmatrix}\bigg{)}.$
@@
for any lift $\textbf{p}_{i}$ of $p_{i}$.
@@
It is straightforward to verify that ${\tt Z}(\tau_{1},\tau_{2})$ is invariant under the action of $\mathrm{PU}(2,1)$.
@@
[Wil12, Proposition 6] Let $\tau_{1}=(p_{1},p_{2},p_{3})$ and $\tau_{2}=(p_{3},p_{4},p_{1})$ be two adjacent real ideal triangles with ${\tt Z}(\tau_{1},\tau_{2})=z=xe^{i\alpha}\in\mathbb{C}\setminus\{-1,0\}$. Then there is an unique isometry $M_{z}\in\mathrm{PU}(2,1)$ such that $M_{z}(p_{1})=p_{3},\ M_{z}(p_{3})=p_{1}$ and $M_{z}(p_{2})=p_{4}$ (see Figure 1).
@@
where $\mathbb{Z}_{2}$ acts on ($(\mathbb{C}\setminus\{-1,0\})^{6g-6+3k}$ by conjugation.
Note that there are exactly $6g-6+3k$ edges in an ideal triangulation on $S_{g,k}$. To define $\psi((\rho,\phi))$, the ${\tt Z}$-invariant is assigned to each edge of $T$, using the pair of adjacent real ideal triangles determined by the image of $\phi$. We refer to [Wil12] for the precise construction and technical details.
@@
This is an elliptic element of order $3$ that cyclically permutes the 3 vertices of the real ideal triangle $\tau_{0}$ in 2.4. Let $\gamma\in\pi_{1}(S_{g,k})$. We fix a lift of the base point. Then $\widetilde{\gamma}$ can be uniquely homotoped to the edges of $\Gamma_{\widetilde{T}}$ in such a way that the $e$-edges and the $t$-edges appear alternatively, beginning with a $t$-edge (see Figure 3).
@@
Let $\widetilde{\gamma}\sim t_{1}*e_{1}*\cdots*t_{r}*e_{r}$. Then $\rho(\gamma)=M(e_{r})M(t_{r})\cdots M(e_{1})M(t_{1})$ (see [Wil12, Proposition 9]).
@@ 3. Proof of the theorem
In this section, we shall prove 1.3. Let $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$ be a given generic representation, and $\gamma\in\pi_{1}(S_{g,k})$. Let us fix an ideal triangulation $T$ on $S_{g,k}$. Let
@@
be a positive expression for some $J,c_{j},X_{j}$ and $\theta_{j}$ as in 3.1. Then
@@
Let $\gamma\in\pi_{1}(S_{g,k})$. Then $\rho(\gamma)\in\mathrm{PU}(2,1)$ has a representative of the form
@@
We notice that the building blocks of $\rho(\gamma)$ are
@@
for some $z_{i}\in\mathcal{Z}$ and $\delta\in\{\pm 1\}$. When $\delta=1$, from equation 3.2, we get
@@
For $\gamma\in\pi_{1}(S_{g,k}),\ \rho(\gamma)$ has a representative in $\mathrm{PU}(2,1)$ such that its trace $tr(\rho(\gamma))$ is a positive expression.
@@
[HJ13, Theorem 8.1.18] Let $A=(a_{ij}),B=(b_{ij})\in M_{n}$ such that $|A|\leq B$, i.e., $|a_{ij}|\leq b_{ij}$ for all $i,j$. Then $\sigma(A)\leq\sigma(|A|)\leq\sigma(B)$, where $|A|=(|a_{ij}|)$ for all $i,j$.
@@
We multiply the first row and first column with $-1$ to get the second and third equality respectively. This proves the lemma. ∎
@@
Now using 3.3, we get
[…]
@@
If $x_{1}x_{2}\cdots x_{r}=1$, then $\ell(\rho(\gamma))=0=\ell(\rho_{0}(\gamma))$. ∎
@@
$|tr(\rho(\gamma))|\leq|tr(\rho_{0}(\gamma))|$
@@
is monotone increasing on $[1,\infty)$ and have a global minima at $x=1$ with $f(1)=2$.
Let $re^{i\phi},r^{-1}e^{i\phi},e^{-2i\phi}$ and $r_{0},r_{0}^{-1},1$ be the eigenvalues of $\rho(\gamma)$ and $\rho_{0}(\gamma)$ respectively. Then
$|tr(\rho(\gamma))|=|re^{i\phi}+r^{-1}e^{i\phi}+e^{-2i\phi}|\leq r+r^{-1}+1\leq r_{0}+r_{0}^{-1}+1=|tr(\rho_{0}(\gamma))|.$
This completes proof. ∎
@@
Note that, 1.5 can be independently proved from 3.5. Let $tr(\rho(\gamma))=p_{t}$ for some positive expression $p_{t}$. Then we can see that $tr(\rho_{0}(\gamma))=p_{t}(0)$. Now from 3.3, we get
$|tr(\rho(\gamma))|=|p_{t}|\leq p_{t}(0)=|tr(\rho_{0}(\gamma))|.$
@@
While domination holds with respect to the Bergman translation length and the modulus of the trace, the discriminator function
@@
which is evaluated at the traces of $\rho(\gamma)$ and $\rho_{0}(\gamma)$ (see 2.1), exhibits more subtle behavior and does not satisfy a similar domination property. To better understand this discrepancy, we conducted numerical experiments using Mathematica. These computations involve evaluating $f(\mathrm{tr}(\rho(\gamma)))$ and $f(\mathrm{tr}(\rho_{0}(\gamma)))$ where $\rho(\gamma)=M_{z_{1}}\mathcal{E}M_{z_{2}}\mathcal{E}^{-1}$ but with randomly sampled parameters $z_{1},z_{2}$ in order to test whether domination holds for the discriminator function.

### v2

@@ 2.3. The $T$-bent representations and their $\tt Z$}-invariants
This part mainly revisits [Wil12] and is included here for the sake of completeness. We begin by recalling a geometric property about pairs of adjacent real ideal triangles. This property is the foundation for defining the $\tt Z$}-invariant.
@@
[…]
@@
for any lift $\textbf{p}_{i}$ of $p_{i}$. Equivalently,
${\tt Z}(\tau_{1},\tau_{2}):=-\frac{\langle\textbf{p}_{4},\textbf{p}_{1}\boxtimes\textbf{p}_{3}\rangle\langle\textbf{p}_{2},\textbf{p}_{1}\rangle}{\langle\textbf{p}_{2},\textbf{p}_{1}\boxtimes\textbf{p}_{3}\rangle\langle\textbf{p}_{4},\textbf{p}_{1}\rangle}.$
@@
It is straightforward to verify that ${\tt Z}(\tau_{1},\tau_{2})$ is invariant under the action of $\mathrm{PU}(2,1)$ and ${\tt Z}(\tau_{1},\tau_{2})=\overline{{\tt Z}(\tau_{2},\tau_{1})}$.
@@
[Wil12, Proposition 6] Let $\tau_{1}=(p_{1},p_{2},p_{3})$ and $\tau_{2}=(p_{3},p_{4},p_{1})$ be two adjacent real ideal triangles with ${\tt Z}(\tau_{1},\tau_{2})=z=xe^{i\alpha}\in\mathbb{C}\setminus\{-1,0\}$. Then there is a unique real symmetry $M_{z}$ such that $M_{z}(p_{1},p_{3})=(p_{3},p_{1})$ and $M_{z}(p_{2},p_{4})=(p_{4},p_{2})$ (see Figure 1).
@@
where $\mathbb{Z}_{2}$ acts on $(\mathbb{C}\setminus\{-1,0\})^{6g-6+3k}$ by conjugation.
Note that there are exactly 6g6+3k6g-6+3k edges in an ideal triangulation on $S_{g,k}$. To define $\psi((\rho,\phi))$, the $\tt Z$}-invariant is assigned to each edge of $T$, using the pair of adjacent real ideal triangles determined by the image of $\phi$. We refer to [Wil12] for the precise construction and technical details.
@@
For the opposite orientation of a $t$-edge, we assign $\mathcal{E}^{-1}$. This is an elliptic element of order $3$ that cyclically permutes the 3 vertices of the real ideal triangle $\tau_{0}$ in 2.4. Let $\gamma\in\pi_{1}(S_{g,k})$. We fix a lift of the base point. Then $\widetilde{\gamma}$ can be uniquely homotoped to the edges of $\Gamma_{\widetilde{T}}$ in such a way that the $e$-edges and the $t$-edges appear alternatively, beginning with a $t$-edge (see Figure 3).
@@
Let $\widetilde{\gamma}\sim t_{1}*e_{1}*\cdots*t_{r}*e_{r}$. If certain $t_{i}$ are oriented in the direction opposite to $\widetilde{\gamma}$, we have $M(t_{i})=\mathcal{E}^{-1}$. Then $\rho(\gamma)$ belongs to the conjugacy class of $M(e_{r})M(t_{r})\cdots M(e_{1})M(t_{1})=M_{z_{r}}\mathcal{E}^{\delta_{r}}\cdots M_{z_{1}}\mathcal{E}^{\delta_{1}}$ where $\delta_{i}=\pm 1$ (see [Wil12, Proposition 9]).
Remark 5.
Note that this process determines $\rho(\gamma)$ only up to conjugacy. Since eigenvalues are invariant under conjugation, determining $\rho(\gamma)$ up to conjugacy is sufficient for our proof. With a slight abuse of notation, we shall therefore write
$\rho(\gamma)=M_{z_{r}}\mathcal{E}^{\delta_{r}}\cdots M_{z_{1}}\mathcal{E}^{\delta_{1}}.$
@@ 3. Proof of the theorem
Notation and Conventions.
Throughout this paper, we view a representation $\rho:\pi_{1}(S_{g,k})\to\mathrm{PU}(2,1)$ as a map into the projective unitary group. Technically, for each $\gamma$, $\rho(\gamma)$ is an equivalence class of matrices in $\mathrm{U}(2,1)$ modulo multiplication by scalars of modulus one. However, whenever convenient we choose a representative $\tilde{\rho}(\gamma)\in\mathrm{U}(2,1)$ and write expressions such as
$\rho(\gamma)=M_{z_{r}}\mathcal{E}^{\delta_{r}}\cdots M_{z_{1}}\mathcal{E}^{\delta_{1}},$
meaning that $\tilde{\rho}(\gamma)=M_{z_{r}}\mathcal{E}^{\delta_{r}}\cdots M_{z_{1}}\mathcal{E}^{\delta_{1}}$ for some choice of lift. All quantities we use (e.g. spectral radius $\sigma$, Bergman translation length $\ell=2\ln\sigma$, and $|\mathrm{tr}|$) are invariant under multiplication by a unit scalar, so this convention does not affect any statement or proof.
We now proceed to the proof of the main theorem. Let $\rho:\pi_{1}(S_{g,k})\rightarrow\mathrm{PU}(2,1)$ be a given generic representation, and $\gamma\in\pi_{1}(S_{g,k})$. Let us fix an ideal triangulation $T$ on $S_{g,k}$. Let
@@
The following proposition is straightforward, and we leave the proof to the reader.
@@
be a positive expression for some $J,c_{j},X_{j}$ and $\theta_{j}$ as in 3.1. Let $p_{t}(0)$ denote the positive expression $p_{t}$ with all its $\theta_{j}$ replaces by 0. Then
@@
Let $\gamma\in\pi_{1}(S_{g,k})$. Then $\rho(\gamma)\in\mathrm{PU}(2,1)$ has a matrix representative of the form
@@
We have seen that the building blocks of $\rho(\gamma)$ (up to conjugacy) are
@@
for some $z_{i}\in\mathcal{Z}$ and $\delta\in\{\pm 1\}$, i.e., powers of all the $\mathcal{E}$ are same (see Figure 4 and [CTT20, Section 5.3]). When $\delta=1$, from equation 3.2, we get
@@
[…]
Figure 4. Segment of a lift of a peripheral loop $\gamma$.
@@
For $\gamma\in\pi_{1}(S_{g,k}),\ \rho(\gamma)$ has a representative in $\mathrm{U}(2,1)$ such that its trace $\mathrm{tr}(\rho(\gamma))$ is a positive expression.
Notation and Conventions.
For two matrices $A=(a_{ij})_{n\times n}$ and $B=(b_{ij})_{n\times n}$, we denote $|A|=(|a_{ij}|)_{n\times n}$, and we write $A\leq B$ if $a_{ij}\leq b_{ij}$ for all $i,j$.
@@
[HJ13, Theorem 8.1.18] Let $A=(a_{ij}),B=(b_{ij})\in M_{n}$ such that $|A|\leq B$. Then $\sigma(A)\leq\sigma(|A|)\leq\sigma(B)$.
@@
We multiply the first row and first column with 1-1 to get the second and third equality respectively. This proves the lemma. ∎
Remark 7.
Note that $A$ and $B$ are conjugate by the matrix $S=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix}$.
@@
Let $B(\gamma)$ denote the matrix
$B(\gamma):=\begin{pmatrix}p_{1}(0)&p_{2}(0)&p_{3}(0)\\ p_{4}(0)&p_{5}(0)&p_{6}(0)\\ p_{7}(0)&p_{8}(0)&p_{9}(0)\end{pmatrix}$.
Then using 3.3, we get
$|\rho(\gamma)|=\begin{pmatrix}|p_{1}|&|-p_{2}|&|-p_{3}|\\ |-p_{4}|&|p_{5}|&|p_{6}|\\ |-p_{7}|&|p_{8}|&|p_{9}|\end{pmatrix}\leq\begin{pmatrix}p_{1}(0)&p_{2}(0)&p_{3}(0)\\ p_{4}(0)&p_{5}(0)&p_{6}(0)\\ p_{7}(0)&p_{8}(0)&p_{9}(0)\end{pmatrix}=B(\gamma).$
@@
If $x_{1}x_{2}\cdots x_{r}=1$, then $\ell(\rho(\gamma))=0=\ell(\rho_{0}(\gamma))$. This completes the proof. ∎
@@
$|\mathrm{tr}(\rho(\gamma))|\leq|\mathrm{tr}(\rho_{0}(\gamma))|$
@@
is monotone increasing on $1,\infty)$ and attains its global minima at $x=1$ with $f(1)=2$.
Let $\{re^{i\phi},r^{-1}e^{i\phi},e^{-2i\phi}\}$ and $\{r_{0},r_{0}^{-1},1\}$ be the eigenvalues of $\rho(\gamma)$ and $\rho_{0}(\gamma)$ respectively. Then
$|\mathrm{tr}(\rho(\gamma))|=|re^{i\phi}+r^{-1}e^{i\phi}+e^{-2i\phi}|\leq r+r^{-1}+1\leq r_{0}+r_{0}^{-1}+1=|\mathrm{tr}(\rho_{0}(\gamma))|.$
This completes the proof. ∎
@@
Note that, 1.6 can be independently proved from 3.5. Let $\mathrm{tr}(\rho(\gamma))=p_{t}$ for some positive expression $p_{t}$. Then we can see that $\mathrm{tr}(\rho_{0}(\gamma))=p_{t}(0)$. Now from 3.3, we get
$|\mathrm{tr}(\rho(\gamma))|=|p_{t}|\leq p_{t}(0)=|\mathrm{tr}(\rho_{0}(\gamma))|.$
@@
While domination holds with respect to the Bergman translation length and the modulus of the trace, it is a natural question whether the domination is preserved under the discriminant function
@@
which is evaluated at the traces of $\rho(\gamma)$ and $\rho_{0}(\gamma)$ (see 2.1). We thank John R. Parker for asking this question. In this Appendix we collect evidences to see that the function $f$ exhibits more subtle behavior and does not satisfy a similar domination property. To better understand this discrepancy, we conducted numerical experiments using Mathematica. These computations involve evaluating $f(\mathrm{tr}(\rho(\gamma)))$ and $f(\mathrm{tr}(\rho_{0}(\gamma)))$ where $\rho(\gamma)=M_{z_{1}}\mathcal{E}M_{z_{2}}\mathcal{E}^{-1}$ but with randomly sampled parameters $z_{1},z_{2}$ in order to test whether domination holds for the discriminator function.